When 519 J up he is added to a sample of liquid water the temperature rises from 21.6 Celsius to 61.8 Celsius given the specific heat capacity of liquid water is 4.184J/G•K how many grams of water are in the sample


1
Expert's answer
2021-03-10T06:01:32-0500

ΔT = 61.8 – 21.6 = 40.2 ºC

Q = mcΔT

"m = \frac{Q}{cΔT} \\

= \frac{519}{4.184 \times 40.2} \\

= \frac{519}{168.19} \\

= 3.08 \;g"

Answer: 3.08 g